# Groups admitting Wirtinger presentations and Gromov hyperbolic groups

### v1

@@ 1. Introduction
Examples of groups that admit Wirtinger presentations include free and free abelian groups, braid groups and pure braid groups, Artin groups, Thompson’s group $F$, and higher-dimensional knot groups. As for twisted Wirtinger presentations, examples include the infinite dihedral group, elementary abelian $2$-groups, Coxeter groups, twisted Artin groups, and cactus groups.
More generally, let $M$ be a closed smooth $n$-dimensional manifold smoothly embedded in $\mathbb{R}^{n+1}$. Then the fundamental group of the complement $\mathbb{R}^{n+1}\setminus M$ admits a twisted Wirtinger presentation, and it admits a Wirtinger presentation if $M$ is orientable [MR1841759].
@@
We refer to [MR672956] for the standard facts about group homology. It is worth noting that any finitely generated abelian group $A$ can be realized as the second homology group $H_{2}(G;\mathbb{Z})$ of a group $G$ admitting a Wirtinger presentation [MR0635591]. Hence, a large number of groups admitting twisted Wirtinger presentations cannot be Gromov hyperbolic. Conversely, many Gromov hyperbolic groups have positive second Betti number, and such groups cannot admit twisted Wirtinger presentations.
@@ 2. Proof
where the right-hand side represents a normalized $2$-cycle of $G$. This homology class is sometimes referred to as the Pontryagin product of $g,h\in G$. It is natural with respect to group homomorphisms. Namely, for a group homomorphism $f\colon G\to H$ and commuting elements $g,h\in G$, one has
@@
Note that the integral homology class $(1,0)\wedge(0,1)\in H_{2}(\mathbb{Z}\times\mathbb{Z})$ can be shown to be a generator of $H_{2}(\mathbb{Z}\times\mathbb{Z})\cong\mathbb{Z}$.
@@
Let $\phi\colon\mathbb{Z}\times\mathbb{Z}\to G$ be a group homomorphism, and let $A\subset G$ be the image of $\phi$. It follows from the conditon (1) that $A$ is an abelian group generated by two elements and of rank at most one. For such an abelian group $A$, the second rational homology group $H_{2}(A;\mathbb{Q})=H_{2}(A)\otimes\mathbb{Q}$ is trivial. This implies that the induced homomorphism $\phi_{*}\colon H_{2}(\mathbb{Z}\times\mathbb{Z};\mathbb{Q})\to H_{2}(G;\mathbb{Q})$ is the zero map, since $\phi_{*}$ factors through $H_{2}(A;\mathbb{Q})$ as
@@
We now assume, in addition, that $G$ is torsion-free. Let $A$ be the kernel of a homomorphism $\phi\colon\mathbb{Z}\times\mathbb{Z}\to G$ as before. Then, by the condition (1), the subgroup $A$ must be either trivial or infinite cyclic. For such a group $A$, the second integral homology $H_{2}(A)$ vanishes. This implies that the induced homomorphism $\phi_{*}\colon H_{2}(\mathbb{Z}\times\mathbb{Z})\to H_{2}(G)$ is the zero map. The second half of the theorem can be proved in a similar manner. ∎

### v2

@@ 1. Introduction
Examples of groups that admit Wirtinger presentations include free and free abelian groups, braid groups and pure braid groups, Artin groups, Thompson’s group $F$, and higher-dimensional knot groups. As for twisted Wirtinger presentations, examples include the infinite dihedral group, elementary abelian 22-groups, Coxeter groups, twisted Artin groups, and cactus groups.
More generally, let $M$ be a closed smooth $n$-dimensional manifold smoothly embedded in $\mathbb{R}^{n+2}$. Then the fundamental group of the complement $\mathbb{R}^{n+2}\setminus M$ admits a twisted Wirtinger presentation, and it admits a Wirtinger presentation if $M$ is orientable [MR1841759].
@@
We refer to [MR672956] for the standard facts about group homology. It is worth noting that any finitely generated abelian group AA can be realized as the second homology group $H_{2}(G;\mathbb{Z})$ of a group $G$ admitting a Wirtinger presentation [MR0635591]. Hence, a large number of groups admitting twisted Wirtinger presentations cannot be Gromov hyperbolic. Conversely, many Gromov hyperbolic groups have positive second Betti number, and such groups cannot admit twisted Wirtinger presentations.
@@ 2. Proof
where the right-hand side represents a normalized 22-cycle of $G$. This homology class is sometimes referred to as the Pontryagin product of $g,h\in G$. It is natural with respect to group homomorphisms. Namely, for a group homomorphism $f\colon G\to H$ and commuting elements $g,h\in G$, one has
@@
Note that the integral homology class $(1,0)\wedge(0,1)\in H_{2}(\mathbb{Z}\times\mathbb{Z})$ can be shown to be a generator of $H_{2}(\mathbb{Z}\times\mathbb{Z})\cong\mathbb{Z}$. For the properties of the Pontryagin product, see [akita-even-Artin].
@@
Let $\phi\colon\mathbb{Z}\times\mathbb{Z}\to G$ be a group homomorphism, and let $A\subset G$ be the image of $\phi$. It follows from the condition (1) that AA is an abelian group generated by two elements and of rank at most one. For such an abelian group AA, the second rational homology group $H_{2}(A;\mathbb{Q})=H_{2}(A)\otimes\mathbb{Q}$ is trivial. This implies that the induced homomorphism $\phi_{*}\colon H_{2}(\mathbb{Z}\times\mathbb{Z};\mathbb{Q})\to H_{2}(G;\mathbb{Q})$ is the zero map, since $\phi_{*}$ factors through $H_{2}(A;\mathbb{Q})$ as
@@
We now assume, in addition, that $G$ is torsion-free. Let AA be the image of a homomorphism $\phi\colon\mathbb{Z}\times\mathbb{Z}\to G$ as before. Then, by the condition (1), the subgroup AA must be either trivial or infinite cyclic. For such a group AA, the second integral homology $H_{2}(A)$ vanishes. This implies that the induced homomorphism $\phi_{*}\colon H_{2}(\mathbb{Z}\times\mathbb{Z})\to H_{2}(G)$ is the zero map. The second half of the theorem can be proved in a similar manner. ∎